OpenStudy (anonymous):

How to find height of a triangle with base of 14 ft.

OpenStudy (ineedhelplz):

Welcome to openstudy! :)

OpenStudy (shadowlegendx):

Is there any addition information that you can provide, like an image or another measurement/dimension?

OpenStudy (welshfella):

you would need to know the area of the triangle as well as the base to find the heigh Area = 1/2 * base * height

OpenStudy (ineedhelplz):

If you use the Pythagorean Theorem for each right triangle and solve each for y2, you can come up with an equation with just x in it and solve for x either by factoring or by using the Quadratic Formula.

OpenStudy (anonymous):

The area is 63 ft square & the base is 14 ft.

OpenStudy (shadowlegendx):

\[63 = \frac{ 1 }{ 2 } \times 14x\]

OpenStudy (shadowlegendx):

Area of a Triangle = 1/2 times the base of the triangle times the height of the triangle

OpenStudy (shadowlegendx):

Solve for x

OpenStudy (shadowlegendx):

Multiply both sides by 2 126 = 14x Divide both sides by 14 x = 9

OpenStudy (shadowlegendx):

The height of your triangle is 9ft

OpenStudy (anonymous):

Thanks so much guys!! :)